answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

Your room size is by measuring by foot 12 inch width and 11 inch length you order queen bed with dresser and armoir it will accommodate or not?

Your units are incorrect. Do you mean feet? Even at 12 X 11 feet, if doors open into the room, it will be a tight fit but it will fit.

How many different ways can you arrange the number 7532?

Alright try to follow my logic so this first digit can either be a 7,5,. hope that helped3, or 2 If seven is the first digit then 5 can be the second, third, or fourth digit and same with the 3 and 2 this gives u six possibilities now multiply by four because 7,5,3,2 can be the first digit and u get 24 possibilities and easier way to do it is 4 factorial (4*3*2*1)

Is the ace a face card?

No. Face cards are jacks, queens, and kings - cards with faces on them.

What is it called when you pick an object and do not return it in probability?

When you pick an object and do not return it, in probability it is termed "without replacement".

What is the probability of having 3 male offspring in a row?

There are factors such as genetic tendencies, age etc which favour male or female offspring so that the probability of a male offspring is not uniformally 1/2.

Next, the question does not specify how many offspring in all, so the possible sequences could be

[any sequence whether or not containing MMM] followed by [MMM] followed by [any sequence whether or not containing MMM].

As the sequence grows longer, the probability that somewhere in the sequence there are 3 male offspring in a row increases.

For 3 offspring, the probability is 1/8 = 0.125

for 8 offspring, the probability is 107/256 = 0.418

The probabilities would need to be weighted together using the proportion of families that have 3 children, 4 children, and so on.

What can be used to make predictions?

A educated guess or hypothesis depending on what it is like something simple (predicting a math or science question).

What is a fair dice?

A fair die is a die where the probability of throwing each of the faces is equal. Dice is two or more die's.

What is a number 0 to1 that describes how likely an event is to occur?

A number between 0 and 1 represents a decimal probability of how likely an event is to occur. It can be converted into a percentage by multiplying it by 100. For instance 0.5x100 = 50% so 0.5 means that something has a 50% chance of occurring. 0 means that the event definitely won't occur, and 1 means that it definitely will.

Rolling an odd number or a 6 on a number cube?

There are 6 numbers total; odd numbers are 1, 3, 5; then a 6 is added so the probability of rolling an odd number or a 6 on a number cube is 4/6 or 2/3 or 66.67%.

What is the probabilty of drawing a 3 from a deck of 52 cards?

There is a 4 in 52 probability (reduce to 1 in 13) of drawing a '3' from a 52-card deck.

How do you write 1234 in expand form?

Expanded Notation of 1,234 = (1 x 103) + (2 x 102) + (3 x 101) + (4 x 100)

What are the odds in favor of getting 2 heads when a coin is tossed twice?

One in four. 1:4.

The probability of getting heads when a fair coin is tossed is: P(H) = 1/2.

The probability of getting heads on a second toss is: P(H) = 1/2, this result is

independent of the result of the first toss.

The probability of having both events happen (heads on the first and heads on the

second toss) is: P(H1UH2) = (1/2)∙(1/2) = 1/4 = 0.25 = 0.25%

Solution to a bag contains 4 balls Two balls are drawn at random and are found to be white What is the probability that all balls are white?

Two balls are certainly white. So the remaining balls may be one of the following
{WW, WN, NW, NN}
where W means a white ball and N means non-white

ball
Let the events be:
E- all four balls are white(WW)


F- 3 balls are white(WN,

NW)
G- only two balls are white(NN)


Now, P(E) =

P(G) =

1/4, and P(F) =

1/2
Let A be the event that the two balls drawn are white
P(A|E)

=

4C2/4C2 =

1
P(A|F)

=

3C2/4C2 =

1/2
P(A|G)

=

2C2/4C2 =

1/6
Now By Bayes' Theorem,
required probability =

P(E|A)

=

(1*1/4) / ( 1*1/4 + 1/2*1/2 + 1/6*1/4 )
=

(1/4) / ( 13/24 )
=

6/13
----------------------------------------------------------------------------------------------------

2nd opinion

Let's say we have 3 boxes with 4 balls each.

Box A has 4 white balls.

Box B has 3 white balls.

Box C has 2 white balls.

The probability of drawing 2 W balls from;

Box A, P(2W│A) =

(4/4)∙(3/3) =

1

Box B, P(2W│B) =

(3/4)∙(2/3) =

1/2

Box C, P(2W│C) =

(2/4)∙(1/3) =

1/6

Say the probability of picking any of the 3 boxes is the same, we have;

P(A) =

1/3

P(B) =

1/3

P(C) =

1/3

Question is, given the event of drawing 2 W balls from a box taken blindly

from the 3 choices, what is the probability that the balls came from box A,

P(A│2W).

Recurring to Bayes Theorem:

P(A│2W) =

[P(A)P(2W│A)]/[P(A)P(2W│A)+P(B)P(2W│B)+P(C)P(2W│C)] =

[(1/3)(1)]/[(1/3)(1)+(1/3)(1/2)+(1/3)(1/6)] =

6/10 =

0.60 =

60%

P(A│2W) =

0.60 =

60%

How is probability usually written?

P(s) = s/t where success = s and total outcomes = t

instead of a fraction it can be written as a decimal.

Two dice tossed Probability sum is 6?

Well both dice have 6 sides for 6 + 6 = 12 and that shall be your denomenator.

The probability that both sides will equal 6 could be..

1,5
2,4
3,3
4,2
5,1

so the answer would be 5/12


* * * * *


Absolute rubbish.

The denominator (not denomenator!) is 6 * 6 = 36 not 12.

So the answer is 5/36.


What is the event if the probability will equal 1?

It is the probability of an event that will definitely happen.

Is a binomial distribution symmetric?

No, in general is not. It is only symmetric if the probability of success in each trial is 0.5

Trending Questions
What does sample mean in an entire population? When two probabilities are multiplied the probability represents a compound event true or false? How many 4 digits numbers can be made using the digits 0 1 2 3 4 5 6 7 8 9 using each digit only once to give the sum 6? When the value of Z -0.84 what is the area probability to the right? If you roll two dice what is the probability of the event that you roll a total of seven or ten on the two dice? What happens to the confidence interval as the standard deviation of a distribution increases? What is the probability of seeing a rainbow tomorrow? Is it more likely that you will roll a 4 if you roll 1 die or that you will roll a sum of 4 if you roll 2 dice? What is a probability based upon an event that has already occurred? Why does a researcher want to go from a normal distribution to a standard normal distributio? If you have a standard deck of 52 cards. You pick one card from deck and then without putting the firsy card back pick a second card. What is the probablity that both cards will be 9s? What is the probability of 7 out of 20 in decimal form? If X is a normal random variable with standard deviation 4.00 and if the probability that X is more than 5.52 is 1271 then what is the mean of X? How many possible outcomes are there when 4 fair coins are tossed at once? What are the chances of rolling a four or higher on two dice with one roll? What is one way in which identical and fraternal similar? Does an outlier have a greater effect on the standard deviation or interquartile range? The difference between the theoretical and empirical probability? What is the probability of not rolling a 4? What do the 4 ace in a card deck stand for?